I was going to check the TF on each stage of PMC's servo.
Unfortunately, I couldn't find the floppy disc drive, so I slide the sliders (gain, RF power) around. When I add more RF power (from 1V to 7V) to 21.5 MHz EOM, the oscilaltion subsides*.
(*5 mins later I came back to turn down the RF power to 1 V again, and the beam was perfectly locked for a few minutes before fluctuated again)
It's not a long term solution, but I note this for further debugging.
One thing about the gain, I see the TF of the whole PMC servo, when I increase the common gain from 0 to ~8 dB, the magnitude of the TF gets higher. If I increase more gain to ~ 10dB or something, the magnitude goes down . So there might be sth wrong about the opamp that controls the gain.